OpenStudy (anonymous):

Help :) Choose the correct simplification of (x^2/4^5)^4 Answers: x^12y^8 x^8/y^9 y^12 x^8/y^20

OpenStudy (hoblos):

is (y) instead of 4 !!!

OpenStudy (anonymous):

HAHAHA yea!!!

OpenStudy (hoblos):

in general (a/b)^2 = a^n/b^n so (x^2/y^5)^4 = (x^2)^4 / (y^5)^4 = x^8 / y^20

OpenStudy (anonymous):

Thank you. :) ill check back to see if your right. :P

OpenStudy (hoblos):

lol 0k check it ;) anyway yw :)